OpenStudy (gabbyalicorn):

A wheel has a diameter of 16 cm. Approximately how far does it travel in 5 revolutions? A. 502.4 cm B. 251.2 cm C. 100.48 cm D. 80 cm

OpenStudy (gabbyalicorn):

@johnweldon1993 i think 80 what im doing : 16 x 5

OpenStudy (johnweldon1993):

When you travel 1 rotation...how far do you go? *hint...you travel in 1 full circle...so you cover how much distance?

OpenStudy (gabbyalicorn):

360

OpenStudy (johnweldon1993):

Okay right...360 degrees...which is also equal to \(\large 2\pi\) right?

OpenStudy (gabbyalicorn):

well when i did 2 x 3.14 it got me to 6.28

OpenStudy (gabbyalicorn):

which is 100.48

OpenStudy (gabbyalicorn):

divide 100.48 by 16

OpenStudy (gabbyalicorn):

oh wait no i was wrong

OpenStudy (gabbyalicorn):

so 100.48 x 5

OpenStudy (gabbyalicorn):

i just realized that

OpenStudy (gabbyalicorn):

which would be 502.4

OpenStudy (gabbyalicorn):

which would be A

OpenStudy (gabbyalicorn):

:D yay!

OpenStudy (unknownunknown):

c/d = pi. c = pi * d = 16 * pi. Times 5 rotations = 16 * pi * 5 = 251.2

OpenStudy (johnweldon1993):

Hang on 1 second...

OpenStudy (johnweldon1993):

Yes >.< that is correct right there! Dumb mistake!

OpenStudy (gabbyalicorn):

oh kk thanks again

OpenStudy (unknownunknown):

Easy to make

OpenStudy (johnweldon1993):

Thank you @unknownunknown it must be too late for me to think!

OpenStudy (unknownunknown):

Haha, don't worry I make worse all the time.